# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v7

@@ 2.2.2 FD Solution: Control Variate
  • As mentioned before, $\hat{V}_{k+1}^{1D}$ can be seen as option value under a 1D simulation process. However, instead of employing the right hand side in Eqn.2.20 directly for least-squares error reduction, Prop.2 allows one to achieve exactly the same effect by just adding the FD ansatz into the projection basis functions. This helps mitigate computational cost as it saves additional tedious Monte Carlo simulation on $\hat{V}^{1D}$, as well as the variance and covariance calculation in Eqn.2.18.
  • From the proof of Prop.2, up to first order least-squares errors, we don’t see benefit of introducing cross term between $\vec{\phi}^{\text{FD}}(x)$ and $\vec{\phi}_{R-1}^{\text{LSM}}(x)$. This simplifies the construction of new basis and avoid overfitting due to further introduction of higher order terms.
  • The most effective FD auxiliary variable $\hat{V}^{1D}$ is obtained by achieving the highest possible correlation with $V$ based on Eqn.2.21. This provides an important theoretical foundation on the choice of FD ansatz in FD-LSM. For instance, under a multi-asset single-factor setting (e.g. Black Scholes or Local Volatility), we can choose the ansatz from single-factor 1D PDE after inter-asset dimension reduction; for single-asset multi-factor settings (e.g. Heston stochastic volatility), one can construct the equivalent 1D process (Local volatility in general) by matching same marginal distributions, see Gyöngy, (1986); finally, for the most general case of multi-asset multi-factor setting, one could apply the dimension reduction on intra-asset level (reducing the factors other than spot, e.g. stochastic volatility on the same asset) to start with, and then on inter-asset level, eventually arriving at 1D problem which is solvable under Eqn.2.16.
@@ 2.2.3 Error Analysis
After establishing the role of FD solution as a control variate in FD-LSM, we proceed with error analysis in the form of Eqn.2.1 as final solution of Least-Squares Monte Carlo.
From this point on, for notational simplicity, we drop the spatial dependence of random variables and condense notation by replacing arguments depending on time by subscripts and further replace $t_{k}$ with simply $k$. For example, $V_{k+1}=V(\omega,\tau(t_{k+1}))$, $Z_{k}=Z(\omega,t_{k})$, $F_{k}=F(\omega,t_{k})$ and $F^{P}_{k}=F^{P}(\omega,t_{k})$. We also drop the $t_{0}$ filtration for total expectation, variance and covariance operators by default unless explicitly specified, i.e, $\operatorname{\mathbb{E}}(\cdot)=\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{0})$, $\operatorname{\mathrm{Var}}(\cdot)=\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{0})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot)=\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{0})$.
First of all, we introduce below Lemma to pin down the local regression error introduced to conditional expectation in Eqn.2.1, i.e., $\sqrt{\operatorname{\mathbb{E}}[(F^{P}_{k}-F_{k})^{2}]}$, under a given projection method $P$ (which can be either $P^{\text{LSM}}$ or $P^{\text{FD-LSM}}$) at $t_{k}$.
@@
Under the same conditions (e.g., $N_{R}$), the local regression error (upper bound) of FD-LSM is reduced, compared to LSM, by a scaling factor $\sqrt{1-\tilde{\rho}^{2}(V_{k+1},\hat{V}_{k+1}^{1D})}$, where $\tilde{\rho}(V_{k+1},\hat{V}_{k+1}^{1D})$ is defined in Eqn.2.22.
@@
From Eqn.4.4 in Appendix A.3.2, one can relate the upper bound of the local regression error, denoted by $\sqrt{\Xi^{P}_{k}[e^{-r\Delta T}V_{k+1}]}$, to the option value variance, after re-labeling $P=P^{\text{LSM}}$, as
$\Xi^{P^{\text{LSM}}}_{k}[e^{-r\Delta T}V_{k+1}]=\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})]$ (2.24)
@@
The result of Eqn.2.24 is consistent with previous studies on the the asymptotic convergence rate of $\Xi^{P^{\text{LSM}}}_{k}$ going to zero under big $O_{p}$ notation by Stentoft, (2001) in their Theorem 1: the dominant term in the asymptotic from is related to the option variance and thus inversely proportional to $N_{R}$; whereas the other term $\sim R^{-\frac{2s}{d}}$ with $s$ denoting the number of continuous derivatives of $F_{k}$ that exist, is negligible in the case of large $s$, i.e., a smooth $F_{k}$ [^5] In the case of WIC in Appendix A.2, there are discontinuities on the digital coupons and the knocked-in short put option, respectively. We assume that such discontinuities can mitigated to ensure smoothness by introducing overhedges, e.g., call spreads, which it is the case in practice., see Newey, (1997).
Combing above with Eqn.2.20 and 2.21, we have
@@
$\begin{split}\Xi_{k}^{P^{\text{FD-LSM}}}[e^{-r\Delta T}V_{k+1}]&=\Xi_{k}^{P^{\text{LSM}}}[e^{-r\Delta T}(V_{k+1}-\beta^{\text{FD}}_{k}\hat{V}^{1D}_{k+1})]\\ &=[1-\tilde{\rho}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\Xi_{k}^{P^{\text{LSM}}}[e^{-r\Delta T}V_{k+1}],\end{split}$ (2.25)
@@
In the light of Lemma 1, we states the main theoretical result in this section as follows.
@@
As a result of aggregation among a generic multi-period (or multi-bucket) setting, the final pricing error of FD-LSM is diminished, compared to LSM, by a scaling factor of $\sqrt{1-\tilde{\rho}^{2}(V,\hat{V}^{1D})}$. Without loss of generality, a flat correlation on option values has been assumed, i.e., $\tilde{\rho}(V_{k+1},\hat{V}_{k+1}^{1D})=\tilde{\rho}(V,\hat{V}^{1D})$ for all $k$.
@@
The proof of Prop.3 will be conducted for American-style option PV in Eqn.2.8 and CVA in Eqn.2.12, separately. As we can see, the derivations for the two categories (with or without the need of dynamic programming) are generic and representative enough to cover all practical use cases of Eqn.2.1.
@@
Here we tackle the error analysis on Eqn.2.8 by considering a most generic multi-period setting, which give rises to the error accumulation under dynamic programming on LSM and FD-LSM. This piece of analysis is largely an extension of previous studies, with a focus on relating the accumulated option pricing errors to local regression error during the backward evolution. We restrict the discussion in the context of holder exercisability, but the conclusion can be applied to issuer exercisability equivalently.
With the discounting factor from period to period $e^{-r\Delta T}$ dropped out for ease of illustration from now on (see Clément et al., (2002)), the dynamic programming solution of option value is re-written here from Eqn.2.7 as
$V_{k}=Z_{k}\Theta(Z_{k}-F_{k})+V_{k+1}[1-\Theta(Z_{k}-F_{k})].$ (2.26)
In a similar way, the approximated solution under least-squares Monte Carlo defined by projection $P$ follows:
$V_{k}^{P}=Z_{k}\Theta(Z_{k}-P_{k}[V_{k+1}^{P}])+V_{k+1}^{P}[1-\Theta(Z_{k}-P_{k}[V_{k+1}^{P}])],$ (2.27)
where $P_{k}$ can be either $P_{k}^{\text{LSM}}$ or $P_{k}^{\text{FD-LSM}}$ as usual, and $P_{k}[V^{P}_{k+1}]$ is the projection on the option value with error accumulated up to the next step due to approximation from backward.
@@
$\begin{split}\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})&=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}([Z_{k}-V_{k+1}][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})]|\operatorname{\mathcal{F}}_{t_{k}}))+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}([Z_{k}-\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})][\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}[(V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})]\\ &=\operatorname{\mathbb{E}}((Z_{k}-F_{k})[\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])+\\ &\qquad\operatorname{\mathbb{E}}(V_{k+1}^{P}-V_{k+1})\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))+\operatorname{\mathrm{Cov}}[(V_{k+1}^{P}-V_{k+1}),\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})],\\ \end{split}$ (2.28)
where $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}^{P}|\operatorname{\mathcal{F}}_{t_{k}}))=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}^{P}|\operatorname{\mathcal{F}}_{t_{k+1}}))$ and $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}}))=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k+1}}))$ have been applied starting with law of total expectation, i.e., $\operatorname{\mathbb{E}}(\cdot)=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{t_{k}}))$.
We can see easily that $\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})$ in Eqn.2.28 is non-positive by induction starting from $\operatorname{\mathbb{E}}(V_{M}^{P})=\operatorname{\mathbb{E}}(V_{M})$. Intuitively, the reason is that the least-squares approximation introduces a sub-optimal exercise strategy giving rise to biased low error for holder exercisable option price. The last covariance term is arguably small which is neglected by previous analysis on the same matter in Clément et al., (2002); here we would quantify it as non-negative in a more precise measure, given the fact that as $V^{P}_{k+1}\to V_{k+1}$, both sides tend to increase under the covariance operator, indicating a non-negative correlation between the two.

### v8

@@ 2.3.2 FD Solution: Control Variate
  • As mentioned before, $\hat{V}_{k+1}^{1D}$ can be seen as option value under a 1D simulation process. However, instead of employing the right-hand side in Eqn.2.32 directly for RSS reduction, Prop.1 allows one to achieve exactly the same effect by just adding the FD ansatz into the projection basis functions. This helps mitigate computational cost as it saves additional tedious Monte Carlo simulation on $\hat{V}^{1D}$, as well as the variance and covariance calculation in Eqn.2.23.
  • From the proof of Prop.1, up to first order of RSS, we don’t see benefit of introducing cross term between $\vec{\phi}^{P^{\text{FD}}}(x)$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$. This simplifies the construction of new basis and avoid overfitting due to further introduction of higher order terms.
  • The most effective FD auxiliary variable $\hat{V}^{1D}$ is obtained by achieving the highest possible correlation with $V$ based on Eqn.2.33. This provides an important theoretical foundation on the choice of FD ansatz in FD-LSM. For instance, under a multi-asset single-factor setting (e.g. Black Scholes or Local Volatility), we can choose the ansatz from single-factor 1D PDE after inter-asset dimension reduction; for single-asset multi-factor settings (e.g. Heston stochastic volatility), one can construct the equivalent 1D process (Local volatility in general) by matching same marginal distributions as mentioned before, see Gyöngy, (1986); finally, for the most general case of multi-asset multi-factor setting, one could apply the dimension reduction on intra-asset level (reducing the factors other than spot, e.g. stochastic volatility on the same asset) to start with, and then on inter-asset level, eventually arriving at 1D problem which is solvable under Eqn.2.20.
@@ 2.3.3 Error Analysis
After establishing the role of FD solution as a control variate in FD-LSM, we proceed with error analysis in the form of Eqn.2.3 as final solution of Least-Squares Monte Carlo.
From this point on, we drop the $t_{0}$ filtration for total expectation, variance and covariance operators by default unless explicitly specified, i.e, $\operatorname{\mathbb{E}}(\cdot)\equiv\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{0})$, $\operatorname{\mathrm{Var}}(\cdot)\equiv\operatorname{\mathrm{Var}}(\cdot|\operatorname{\mathcal{F}}_{0})$ and $\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot)\equiv\operatorname{\mathrm{Cov}}(\cdot,\cdot\cdot|\operatorname{\mathcal{F}}_{0})$.
To initiate the final pricing error analysis, we introduce below lemma to pin down the local regression error, i.e., the mean squared error (MSE) $\sqrt{\operatorname{\mathbb{E}}[(F^{P}_{k}-F_{k})^{2}]}$ of $F^{P}_{k}$ (introduced to conditional expectation in Eqn.2.3) under a given projection method $P$ (either $P^{\text{LSM}}$ or $P^{\text{FD-LSM}}$) at $t_{k}$.
@@
Under the ceteris paribus assumption with $R\geq 3$, the local regression error (upper bound) of FD-LSM is reduced, compared to LSM, by a scaling factor
$\sqrt{\frac{R-1}{R}(1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}_{k+1}^{1D}))}$, where $\tilde{\rho}_{k}(V_{k+1},\hat{V}_{k+1}^{1D})$ is defined in Eqn.2.34.
@@
We denote by $\Xi^{P^{\text{LSM}}}k[e^{-r\Delta T}V{k+1}]$ the upper bound of $\sqrt{\operatorname{\mathbb{E}}[(F^{P^{\text{LSM}}}_{k}-F_{k})^{2}]}$. A detailed derivation of this bound is provided in Appendix A.4. One of the key points is to express $F_{k}$ in terms of exact regression coefficients $\vec{\beta}^{P^{\text{LSM}}0}$ satisfying
$F_{k}=\vec{\beta}^{P^{\text{LSM}}0}_{k}\cdot\vec{\phi}^{P^{\text{LSM}}}_{R,k}.$ (2.36)
From Eqn.4.3 in Appendix, it concludes that
$\Xi^{P^{\text{LSM}}}_{k}[e^{-r\Delta T}V_{k+1}]=R\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})].$ (2.37)
@@
This result is consistent with previous studies on the asymptotic convergence rate of $\Xi^{P^{\text{LSM}}}_{k}$ going to zero under big $O_{p}$ notation by Stentoft, (2001) in their Theorem 1: the dominant term in the asymptotic from is related to the option variance and thus inversely proportional to $N_{R}$; whereas the other term $\sim R^{-\frac{2s}{d}}$ (with $s$ denoting the number of continuous derivatives of $F_{k}$ that exist) is negligible, i.e.,
$R^{-\frac{2s}{d}}\ll 1,$ (2.38)
in the case of $R\geq 2$ and large $s\gg\frac{d}{2}$, i.e., a smooth $F_{k}$ [^6] In the case of WIC in Appendix A.2, there are discontinuities on the digital coupons and the knocked-in short put option, respectively. We assume that such discontinuities can be mitigated to ensure smoothness by introducing overhedges, e.g., call spreads, which aligns with common practice., see Newey, (1997). In fact, this condition is a justification of the exact decomposition in Eqn.2.36.
Apparently, the upper bound estimate using Eqn.2.37 is also valid for FD-LSM considering all the basis functions as a whole, but it is a loose one which ignores the correlation between $\vec{\phi}^{P^{\text{FD-LSM}}}_{R,k}$ and $V_{k+1}$. Thanks to the decomposition in Eqn.2.32 from Prop.1, we can achieve a tighter estimate as follows.
By construction, the MSE of $F^{P^{\text{FD-LSM}}}_{k}$ can be decomposed into individual $L^{2}$-norm square of orthogonal components $P_{k}^{\text{FD}}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}]$ and $M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}]$, respectively.
By definition, $P_{k}^{\text{FD}}$ projects a variable onto $\text{span}(\vec{\phi}^{P^{\text{FD}}}_{k})$ as a subset of $\text{span}(\vec{\phi}^{P^{\text{FD-LSM}}}_{k})$. Therefore, Law of Iterated Projection gives $P_{k}^{\text{FD}}[F^{P^{\text{FD-LSM}}}_{k}]=P_{k}^{\text{FD}}[P^{\text{FD-LSM}}_{k}[e^{-r\Delta T}V_{k+1}]]=P_{k}^{\text{FD}}[e^{-r\Delta T}V_{k+1}]$. Additionally, taking into account Eqn.2.24, one arrives at $P_{k}^{\text{FD}}[F^{P^{\text{FD-LSM}}}_{k}]=P_{k}^{\text{FD}}[F_{k}]$; therefore
$\begin{split}\operatorname{\mathbb{E}}[(F^{P^{\text{FD-LSM}}}_{k}-F_{k})^{2}]&=\operatorname{\mathbb{E}}[(P^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]+\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}-F_{k}])^{2}]\\ &=\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}].\end{split}$ (2.39)
Applying $M^{\text{FD}}_{k}$ to the right-hand side of Eqn.2.32 and exploiting the orthogonality condition $M^{\text{FD}}_{k}[F^{\text{FD}}_{k}[\cdot]]=0$, and then expressing $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ via Eqn.2.25 and Eqn.2.27, we have
@@
$\begin{split}M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]&=M^{\text{FD}}_{k}[P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})]]\\ &=\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}}_{k}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D}))\cdot M^{\text{FD}}_{k}[\vec{\phi}^{P^{\text{LSM}}}_{R-1,k}]\\ &=(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}T}_{R-1,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})]\cdot\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}.\\ \end{split}$ (2.40)
Similar to Eqn.2.36, exact coefficients $\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}0}_{k}$ satisfying $M^{\text{FD}}[F_{k}]=\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}0}_{k}\cdot\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}$ are introduced. According to Eqn.2.27, we have
$\begin{split}\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}0}_{k}&=\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}}_{k}(P_{k}^{\text{FD}}[F_{k}])\\ &=(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}T}_{R-1,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}(F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k}))].\end{split}$ (2.41)
In Eqn.2.40 and 2.41, we have ignored the constant term $\alpha_{k}^{\text{FD}}(F_{k})$ that can be absorbed in the projection $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ as usual.
As discussed before, $R\geq 2$ is required due to the constraint in Eqn.2.38. In fact, to account for the decrement in the number of basis functions when using $\vec{\phi}^{P^{\text{LSM}\perp\text{FD}}}_{R-1,k}$, we would enforce $R\geq 3$. Fortunately, the control variate makes the smoothness condition on $F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})$ easier to achieve than on plain $F_{k}$.
In what follows, the calculation of $\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}]$ mirrors the procedure used for $\operatorname{\mathbb{E}}[(F^{P^{\text{LSM}}}_{k}-F_{k})^{2}]$ after performing below transformation:
$\begin{split}F_{k}&\rightarrow F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})\\ \vec{\phi}^{P^{\text{LSM}}}_{R,k}&\rightarrow\vec{\phi}^{P^{\text{LSM}\perp\text{LSM}}}_{R-1,k}\\ \vec{\beta}^{P^{\text{LSM}}}_{k}(\cdot)&\rightarrow\vec{\beta}^{P^{\text{LSM}\perp{\text{FD}}}}_{k}(\cdot)\\ e^{-r\Delta T}V_{k+1}&\rightarrow e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})\end{split}.$ (2.42)
Along with Eqn.2.39, it yields
$\begin{split}&\operatorname{\mathbb{E}}[(M^{\text{FD}}_{k}[F^{P^{\text{FD-LSM}}}_{k}]-M^{\text{FD}}_{k}[F_{k}])^{2}]\\ \leq&(R-1)\cdot\operatorname{\mathbb{E}}[(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})-(F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})))^{2}]\\ \leq&(R-1)\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}})]\\ =&(R-1)[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})],\end{split}$ (2.43)
where the identity $\operatorname{\mathbb{E}}[e^{-r\Delta T}(V_{k+1}-\beta_{k}^{\text{FD}}(F_{k})\hat{V}_{k+1}^{1D})|\operatorname{\mathcal{F}}_{t_{k}}]=F_{k}-\beta_{k}^{\text{FD}}(F_{k})f^{\text{FD}}_{k}(X_{k})$ and Eqn.2.33 have been used in the penultimate and last line, respectively, apart from following the derivation described in Appendix A.4 with the mapping in Eqn.2.42.
Denoting $\Xi_{k}^{P^{\text{FD-LSM}}}[e^{-r\Delta T}V_{k+1}]$ as the upper bound of $\operatorname{\mathbb{E}}[(F^{P^{\text{FD-LSM}}}_{k}-F_{k})^{2}]$, and comparing above with Eqn.2.37, we have
$\begin{split}\Xi_{k}^{P^{\text{FD-LSM}}}[e^{-r\Delta T}V_{k+1}]=\frac{R-1}{R}[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\cdot\Xi_{k}^{P^{\text{LSM}}}[e^{-r\Delta T}V_{k+1}],\end{split}$ (2.44)
@@
Let’s delve deeper into Lemma 3 before proceeding:
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 1$, the result degenerates to Lemma 1 as expected.
  • In the limit of $\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})\to 0$, the control variate has null effect as $f^{\text{FD}}_{k}(X_{k})$ is orthogonal to $F_{k}$ in $L^{2}$. In this regime, the remaining mild error reduction, going as $\sqrt{\frac{R}{R-1}}$ is purely from reducing susceptibility of overfitting by removing unnecessary high order monomial component under the constraint of Eqn.2.38; further taking the limit of $R\to\infty$, $P_{k}^{\text{LSM},\text{LSM}\perp\text{FD}}$ converges to $P_{k}^{\text{LSM}}$ with MSE being equal to each other.
In the light of Lemma 3 as an excellent implication of Prop.1, we state the main theoretical result, and its assumption(s), in this section as follows.
Assumption 1.
In the option PV problem, below inequality (with $e^{-r\Delta T}$ dropped out) holds for all $k$:
$\operatorname{\mathrm{Var}}(B^{P}_{k})\geq\operatorname{\mathrm{Var}}(V^{P}_{k+1}-V_{k+1})\cdot\mathcal{X}^{P}_{k},$
where
$\begin{split}B^{P}_{k}&:=(Z_{k}-F_{k})(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})),\\ \mathcal{X}^{P}_{k}&:=\operatorname{\mathbb{E}}(\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|)),\end{split}$ (2.45)
with $V^{P}_{k+1}$ denoting the approximated option price under projection $P$, as opposed to the exact $V_{k+1}$.
Assumption 2.
Further defining below quantity in the option PV problem
$\begin{split}\mathcal{Y}^{P}_{k}&:=\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})),\end{split}$ (2.46)
below inequality is true
$\mathcal{X}^{P}_{0}\leq(1-\mathcal{Y}^{P}_{0})^{2},$
in the limit of $M\to\infty$, where $\mathcal{X}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{X}^{P}_{k}$ and $\mathcal{Y}^{P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\mathcal{Y}^{P}_{k}$. For avoidance of doubt, we use “0” as the subscript to represent this limiting case where $k$ is indexed from 1 as $k\in[1,M]$.
@@
The pricing error of FD-LSM converges to zero as $\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})\to 1$, where $m^{*}:=\arg_{m\in[1,M]}\min\tilde{\rho}m(V{m+1},\hat{V}^{1D}_{m+1})$. This convergence holds in a generic multi-period (or multi-bucket) setting, while the error of the LSM counterpart remains finite. For the option PV problem, Assumption 1 is required; furthermore, if Assumption 2 is also satisfied, the convergence of the FD-LSM error remains valid even as $M\to\infty$.
@@
The proof will be conducted for American-style option PV in Eqn.2.12 and CVA in Eqn.2.16, separately. As we can see, the derivations for the two categories (with or without the need of dynamic programming) are generic and representative enough to cover practical use cases of both categories.
@@
Here we tackle the error analysis on Eqn.2.12 by considering a most generic multi-period setting, which give rises to the error accumulation under dynamic programming on LSM and FD-LSM. This piece of analysis is largely an extension of previous studies, with a focus on relating the accumulated option pricing errors to local regression error during the backward evolution. We restrict the discussion in the context of holder exercisability, but the conclusion can be applied to issuer exercisability equivalently.
With the discounting factor from period to period $e^{-r\Delta T}$ dropped out for ease of illustration from now on (see Clément et al., (2002)), the dynamic programming solution of option value for both exact $V_{k}$ and approximated $V_{k}^{P}$ can be re-written here from Eqn.2.11 as
$V_{k}$ $=Z_{k}\Theta(Z_{k}-F_{k})+V_{k+1}(1-\Theta(Z_{k}-F_{k})),$ (2.47)
                        and
$V_{k}^{P}$ $=Z_{k}\Theta(Z_{k}-P_{k}[V_{k+1}^{P}])+V_{k+1}^{P}(1-\Theta(Z_{k}-P_{k}[V_{k+1}^{P}])),$
respectively, where $P_{k}$ can be either $P_{k}^{\text{LSM}}$ or $P_{k}^{\text{FD-LSM}}$ as usual, and $P_{k}[V^{P}_{k+1}]$ is the projection on the option value with error accumulated up to the next step due to approximation from backward. The option PV at $t=0$ is $V_{k=1}$( and its proxy $V^{P}_{k=1}$) due to omitting $e^{-r\Delta T}$.
@@
$\begin{split}\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})&=\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}((Z_{k}-V_{k+1})(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k}))|\operatorname{\mathcal{F}}_{t_{k}}))+\\ &\qquad\operatorname{\mathbb{E}}((V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))\\ &=\operatorname{\mathbb{E}}((Z_{k}-\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}}))(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})))+\\ &\qquad\operatorname{\mathbb{E}}((V_{k+1}^{P}-V_{k+1})\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))\\ &=\operatorname{\mathbb{E}}((Z_{k}-F_{k})(\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})))+\\ &\qquad\operatorname{\mathbb{E}}(V_{k+1}^{P}-V_{k+1})\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))+\operatorname{\mathrm{Cov}}(V_{k+1}^{P}-V_{k+1},\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})),\\ \end{split}$ (2.48)
where $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}^{P}|\operatorname{\mathcal{F}}_{t_{k}}))\equiv\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}^{P}|\operatorname{\mathcal{F}}_{t_{k+1}}))$ and $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}}))\equiv\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(V_{k+1}|\operatorname{\mathcal{F}}_{t_{k+1}}))$ have been applied starting with law of total expectation, i.e., $\operatorname{\mathbb{E}}(\cdot)\equiv\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(\cdot|\operatorname{\mathcal{F}}_{t_{k}}))$.
We can see easily that $\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})$ in Eqn.2.48 is non-positive by induction starting from $\operatorname{\mathbb{E}}(V_{M}^{P})=\operatorname{\mathbb{E}}(V_{M})$. Intuitively, the reason is that the least-squares approximation introduces a sub-optimal exercise strategy giving rise to biased low error for holder exercisable option price. The last covariance term is arguably small which is neglected by previous analysis on the same matter in Clément et al., (2002); here we would quantify it as non-negative in a more precise measure, given the fact that as $V^{P}_{k+1}\to V_{k+1}$, both sides tend to increase under the covariance operator, indicating a non-negative correlation between the two.